OpenStudy (anonymous):

Which number lies between 4.3 and 4.5 on the number line? How to find that answer?

OpenStudy (anonymous):

theres an infinite amount of numbers, it could be 4.4, 4.44444444444444, 4.4999 etc

OpenStudy (anonymous):

if you mean the middle then just take the average (4.3+4.5)/2 = 4.4

OpenStudy (anonymous):

Thanks :)

OpenStudy (anonymous):

Wait do which one of these would be correct? 4 1/5 4 7/20 4.06

OpenStudy (anonymous):

so*

OpenStudy (anonymous):

@jayz657

OpenStudy (anonymous):

just find the decimals for it so 1/5 = 0.2 7/20 - 0.35 4.06 is 4.06 now look at the number line none of these are in between 4.3 and 4.5

OpenStudy (anonymous):

wait dont really know what are those 4s in front of the fractions mean?

OpenStudy (anonymous):

4 and 1/5th and 4 and 7/20th's

OpenStudy (anonymous):

oh right hahah sorry i just noticed let me do it again 4 and 1/5 = 2.5 4 and 7/20 = 4.35 we can see 4.35 is in between 4.3 and 4.5 so thats your answer

OpenStudy (anonymous):

Haha. Thank you! :)

OpenStudy (anonymous):

np glad to help